# ToolSandbox: A Stateful, Conversational, Interactive Evaluation Benchmark for LLM Tool Use Capabilities

### v1

@@ 1 Introduction
Figure 1: An example evaluation trajectory from ToolSandbox. Some message contents and milestones were truncated and streamlined for visual clarity. The Message Bus represents a full dialog history between the User, the Agent and the Execution Environment. The World State represents mutable database snapshots at a given turn. The Milestones represent predefined key events that need to happen in this trajectory. In this example, the User intended to send a message, while cellular service is turned off. The Agent should first understand the User’s intent, and prompt for necessary arguments from the User. After collecting all arguments with the help of the search_contacts tool, the Agent attempted to send the message, figured out it needs to enable cellular service upon failure, and retried. To evaluate this trajectory, we find the best match for all Milestones against Message Bus and World State in each turn while maintaining topological order.
ToolSandbox BFCL ToolEval API-Bank
State Dependency
Conversational
Interactive
Human Authored Ground Truth
Table 1: A comparison between ToolSandbox and other tool-use Benchmarks.
@@
Despite the paradigm shift towards a more simplified problem formulation, the stateful, conversational and interactive nature of task oriented dialog remains, and poses a significant challenge for systematic and accurate evaluation of tool-using LLMs. Existing benchmarks like the Berkeley Function Calling Leaderboard (BFCL) Yan et al. (2024), ToolEval Qin et al. (2024) and API-Bank Li et al. (2023) attempted to tackle some of these challenges, but there is yet to be an all encompassing solution.
@@
BFCL Yan et al. (2024) and ToolEval Qin et al. (2024) both rely on stateless tools interacting with web services (through RESTful APIs). As such, these evaluation benchmarks are designed to assess how agents make trials with a static environment. API-Bank Li et al. (2023) does include a set of tools to modify world states, but does not study the impact of state dependencies.
@@
Conversational evaluation is crucial yet challenging when assessing a dialog policy, due to the interdependency between a user and said policy, as well as the ambiguous nature of natural language. To facilitate automated conversational evaluation, a common practice is to implement a simulated user Zhang et al. (2024); Sekulic et al. (2024). However, BFCL and ToolEval only evaluate self-contained, unambiguous single-turn user queries, which is hardly realistic. API-Bank evaluates on unrolled predefined off-policy dialog trajectories, and thus cannot assess the agent’s performance based on its own policy.
@@
Real world scenarios are full of surprises. The agent could issue an erroneous tool call. Tool execution could raise an unexpected exception. And the user could issue a follow-up correcting a previous statement. An interactive evaluation framework assessing the immediate return of key interactions with user or environment would be necessary to capture the intricate interaction between different roles. Such an interactive evaluation should provide full spectrum and fine-grained evaluation of any multi-turn session. In this regard, BFCL and API-Bank both rely on a predefined trajectory, and by extension rely on static turn wise evaluation metrics. Even though ToolEval allows multiple rounds of interaction between the Agent and tools, it relies solely on an LLM evaluator to judge the final pass rate and win rate of trajectories, which raises questions to its reliability and interpretability.
@@ 2 ToolSandbox Design
At its core, ToolSandbox is a Python native LLM testing environment, with Execution Context as world state abstraction and Python functions as Tools, where User, Agent and Execution Environment communicate with each other through a Message Bus to complete a task, which is evaluated against predefined Milestones and Minefields. As shown in Figure 2, a typical test case starts with the User speaking to the Agent. From then on, the role being addressed gets to speak next, until the end state is reached. Upon receiving a User request, an Agent can decide to respond to the User asking for more information, or inform the Execution Environment to execute a Tool, providing intended tool name and arguments. The Execution Environment executes the Tool in an code.InteractiveConsole, Foundation (2024), which depending on the Tool modifies the world state stored in the Execution Context, and responds to the Agent. Once the User decides the task has been completed, it informs the Execution Environment to execute the end_conversation tool, which puts the system in the end state, ready to be evaluated based on the dialog’s similarity to Milestones and Minefields. The remainder of this section introduces the functionality of each component in more details.
2.1 Tools
Tools in ToolSandbox are a set of highly composable, explicitly or implicitly dependent Python functions, creating complex reasoning challenges. Besides python native tools, a handful of carefully selected RapidAPI tools were also included with a thin layer of python wrapper. Tools manipulate world state through the Execution Context when necessary, and raise informative exceptions when execution conditions were not met. See Appendix A.1 and A.2 for more information. As an example, in Figure 1, when send_message tool is called while cellular service is off, a ConnectionError is raised. This allows the Agent to reason over possible exceptions, and deduce the tool needed to resolve the exception.
To support ablation studies on the effect of tool schema representation on agent accuracy, we have implemented multiple tool augmentations, e.g.:
  • The agent is given distraction tools not needed to complete the task.
  • Tool or argument names becomes less informative, e.g. using a tool name of settings_0 instead of set_cellular_service_status to test if the agent relies on them to infer a tool’s purpose.
  • Information like argument descriptions or type hints are removed.
For more details on the augmentations please refer to Appendix A.2.1.
2.2 Roles and Message Bus
In ToolSandbox there are three roles: User, Agent (Assistant) and Execution Environment. The Execution Environment, as a dedicated role, is responsible for executing tool-use requests from the Agent and returning the results. Interaction between the roles is enabled through a message passing system. Each message contains a sender role, recipient role, content as well as to which roles the message is visible to. A simple orchestrator determines message passing order by allowing the most recent recipient to be the next sender. Instead of representing the conversation as a single message thread, we use a collection of messages, i.e. Message Bus, stored within the Execution Context. The Message Bus contains a linear history of message transactions between all roles. As is shown in Appendix A.3, each role writes to the same Message Bus. However, when reading from the Message Bus, each role can only access a sub-view of the Message Bus based on which roles are allowed to "see" the individual messages. We will introduce each role in the following paragraphs.
User Role
@@ 2.2 Roles and Message Bus
The User role represents a human interacting with an Agent, hoping to complete a task through possibly multiple rounds of conversation. When the User role decides the task has been completed, or could not be completed, it can terminate the conversation using the end_conversation tool, which is the single tool available to the User. The User role is implemented with an LLM (GPT-4o) and carefully calibrated prompting design to make the user simulator more realistic. As related studies in user simulation Zhang et al. (2024); Sekulic et al. (2024) suggest, one should include the user’s overall goal in the simulator’s system prompt. However, we found this is often insufficient for the complex interactions in ToolSandbox, and can lead to two categories of failures. In some cases, it is infeasible for an LLM simulated user to judge task completion, or provide follow-up information with only access to the user goal, and not the expected result, which could lead to hallucination. Also, with only a single system prompt, the simulated user could be derailed by the tool-use agent, failing to follow instructions. Examples of these failures can be found in Appendix A.4.
@@
In light of this, we propose two additional components in user simulator prompts: Knowledge Boundary, which inform the user simulator what it should and should not know, providing partial access to expected result, combating hallucination. And Demonstration, which provides few shot example dialogs to the user simulator. Prompt examples can be found in Appendix A.4. Note that demonstration is only visible to the user simulator and not the agent. We performed an ablation study for these components in Table 2. With both approaches combined, the LLM simulated user achieves the lowest error rate in all categories. User simulator error rate is also found to be consistent across well performing agents, shown in Table 5.
@@
Agent Role
Initially, the Agent role will receive a message from the User in natural language. The Agent could decide to prompt the User again for additional information, or decide to issue a tool call towards the Execution Environment. When issuing a tool call, the Agent selects the name of the tool from a list of available tools and provides necessary arguments, commonly expressed as JSON objects. These JSON objects are converted to executable Python code, see Appendix A.5, and sent to the Execution Environment for execution.
Execution Environment Role
Some LLMs support parallel tool calling, intended to increase efficiency when multiple, independent tools need to be called. However, if an LLM uses parallel tool calls for dependent tools, it should be penalized accordingly. For example, in Figure 1 where the agent has to enable cellular service before sending a text message, parallel tool calls should not be used. Execution Environment handles race conditions in parallel tool calls by following Murphy’s Law, ensuring race condition always happens if detected.
2.3 Evaluation
With an interactive, stateful and conversational environment, evaluation trajectories are highly dynamic. Multiple trajectories can lead to the same outcome. A given task may be completed using different tools, the same tools in a different order, or through trial and error, and the evaluation strategy has to be flexible enough to accommodate for that. To combat this, we developed an evaluation strategy based on Milestones and Minefields, which defines key events that must or must not happen in a trajectory, allowing us to evaluate any trajectory with rich intermediate and final execution signals, providing deeper understanding of the model performance. An example can be found in Figure 10.
@@ 2.3 Evaluation
Each milestone also defines a similarity measure which calculates a 0 to 1 similarity between each turn and the milestone. Types of available similarity measures are introduced in Appendix A.6. Milestones form a directed acyclic graph (DAG) based on temporal dependency. To evaluate a trajectory against a milestone DAG, we find the the highest averaged similarity $\text{score}_{M+}$ among all possible mappings between turns and milestones, given that the resulting chronic milestone sequence is a topological sort of the DAG. Task efficiency is not considered by Milestones, and is instead tracked by a complementary turn count metric shown in Appendix D.2. We introduce the milestone matching process with more details in Appendix A.6.
@@ 3 Test Scenarios
ToolSandbox BFCL ToolEval API-Bank
Avg Turns 13.9 2.00 7.53 3.88
Avg Tool calls 3.80 0.78 1.46 2.04
Test cases 1032 2000 1625 261
Tools 34 1193 3917 73
Table 3: Statistics between ToolSandbox and other tool-use benchmarks. Calculation details can be found in Appendix B.1.
ToolSandbox contains 1032 test cases meticulously crafted by 2 internal domain experts to capture challenging tool-use scenarios, with human authored and carefully calibrated Milestones and Minefields to support evaluation. 1 annotator is tasked to create test scenarios, while the other acts as an agent to validate milestones and minefields. We designed a rigorous annotation process to ensure coverage across realistic, complex use case scenarios, detailed in Appendix B.2. Statistics comparison between ToolSandbox and other benchmarks can be found in Table 3. Tools in ToolSandbox are designed to be representative, diverse and composable in conversational dialogs, while making tool count manageable for milestone annotation. As a result, ToolSandbox test scenarios contain on average much higher number of tool calls and turns per dialog compared to other benchmarks. Additional details about tool domain coverage and design principles can be found in Appendix B.3.
To closely inspect the intricate challenges in LLM tool-use applications, test scenarios are organized into detailed categories, statistics can be found in Appendix B.4. A test scenario is defined by the initial world state, the initial messages, the available tools and the evaluation milestone and minefields. Variations of test scenarios are also introduced using the tool augmentations described in 2.1. Scenarios are organized by the following categories:
@@
These categories apply to scenarios where one / multiple tool calls are needed to fulfill the user task. Examples are shown in Appendix C.2. Note that this definition is different from the Berkeley Function-Calling leaderboard Yan et al. (2024), which resembles distraction tools in ToolSandbox described in Section 2.1.
@@ 4 Evaluation Results
Table 4 shows the average similarity for each of the scenario categories described in Section 3 and tool augmentations described in Section 2.1. There is a significant performance gap between proprietary and open source models, with the best performing open source model Hermes interstellarninja et al. lagging more than 20 points behind the second to last proprietary model Claude-3-Haiku Anthropic (2024). This is partly due to the fact that models like Gorilla Patil et al. (2023) and Command-R Cohere and for AI (2024) are incapable of consuming tool responses, as shown in Appendix D.1. They can theoretically solve Single Tool Call test scenarios, but would fail in any scenario that requires multiple tool calls. As for Hermes and Mistral Jiang et al. (2023), both models struggle at identifying when a tool call should be issued. Mistral for example would often mistake a tool-use scenario for a code generation task, as shown in Figure 11. These models’ subpar performance unexpectedly caused them to achieve higher rating in the Insufficient Information category, which rewards the model for not generating hallucinated tool calls or arguments when provided tools are insufficient to complete the task. This should be considered a side effect instead of a positive outcome.
@@
The State Dependency category shows an interesting trend where, larger models like GPT-4 Achiam et al. (2023) and Claude-3-Opus perform significantly worse than mid to smaller sized models like GPT-3.5-Turbo and Claude-3-Sonnet. This is due to erroneous parallel tool calls in face of state dependency. As mentioned in Section 2.2, the Execution Environment always surfaces race conditions when present. Larger models like GPT-4 and Claude-3-Opus are prune to issuing parallel tool calls even for dependent tools, leading to a performance deficiency. An example is shown in Figure 17. Nested state dependency is also tricky to solve efficiently. As shown in Figure 18, models often forget about open issues and would not optimally backtrack, leading to repeated errors and as a result a much higher than optimal turn count.
@@
Robustness against tool augmentations seems to vary model by model. While adding distraction tools, Claude-3-Sonnet seems to be affected the most, dropping almost 10 points between 0 distraction tools, and making all ToolSandbox tools available. GPT-4o is particularly susceptible to Tool Description Scrambling, GPT-4 pays extra attention to argument descriptions, and Gemini-1.5 doesn’t do well with Argument Type Scrambling.

### v2

@@ 1 Introduction
Figure 1: An example evaluation trajectory from ToolSandbox. Some message contents and milestones were truncated and streamlined for visual clarity. The Message Bus represents a full dialog history between the User, the Agent and the Execution Environment. The World State represents mutable database snapshots at a given turn. The Milestones represent predefined key events that need to happen in this trajectory. The light blue boxes represent the human authored scenario definition, including the user goal, initial user utterance, initial world state and milestones. The light yellow boxes represent on policy rollout collected during an interactive evaluation run. In this example, the User intended to send a message, while cellular service is turned off. The Agent should first understand the User’s intent, and prompt for necessary arguments from the User. After collecting all arguments with the help of the search_contacts tool, the Agent attempted to send the message, figured out it needs to enable cellular service upon failure, and retried. To evaluate this trajectory, we find the best match for all Milestones against Message Bus and World State in each turn while maintaining topological order.
State Dependency Conversational Interactive Human Authored Ground Truth
ToolSandbox
BFCL
ToolEval
API-Bank
ToolTalk
$\tau$-bench
Table 1: A comparison of ToolSandbox and other tool-use benchmarks.
@@
Despite the paradigm shift towards a more simplified problem formulation, the stateful, conversational and interactive nature of task oriented dialog remains, and poses a significant challenge for systematic and accurate evaluation of tool-using LLMs. Existing benchmarks like the Berkeley Function Calling Leaderboard (BFCL) Yan et al. (2024), ToolEval Qin et al. (2024), API-Bank Li et al. (2023), ToolTalk Farn and Shin (2023) and $\tau$-bench Yao et al. (2024) attempted to tackle some of these challenges, but there is yet to be an all encompassing solution.
@@
BFCL Yan et al. (2024) and ToolEval Qin et al. (2024) both rely on stateless tools interacting with web services (through RESTful APIs). As such, these evaluation benchmarks are designed to assess how agents make trials with a static environment. API-Bank Li et al. (2023), ToolTalk Farn and Shin (2023) and $\tau$-bench Yao et al. (2024) does include a set of tools to modify world states, but does not study the impact of state dependencies.
@@
Conversational evaluation is crucial yet challenging when assessing a dialog policy, due to the interdependency between a user and said policy, as well as the ambiguous nature of natural language. To facilitate automated conversational evaluation, a common practice is to implement a simulated user Zhang et al. (2024); Sekulic et al. (2024). However, BFCL and ToolEval only evaluate self-contained, unambiguous single-turn user queries, which is hardly realistic. API-Bank and ToolTalk evaluates on unrolled predefined off-policy dialog trajectories, and thus cannot assess the agent’s performance based on its own policy.
@@
Real world scenarios are full of surprises. The agent could issue an erroneous tool call. Tool execution could raise an unexpected exception. And the user could issue a follow-up correcting a previous statement. An interactive evaluation framework assessing the immediate return of key interactions with user or environment would be necessary to capture the intricate interaction between different roles. Such an interactive evaluation should provide full spectrum and fine-grained evaluation of any multi-turn session. In this regard, BFCL, API-Bank and ToolTalk rely on a predefined trajectory, and by extension relying on static turn wise evaluation metrics. $\tau$-bench requires agent action to match a single predetermined sequence, leaving no room for error correction in the model, or multiple possible answer sequences. Even though ToolEval allows multiple rounds of interaction between the Agent and tools, it relies solely on an LLM evaluator to judge the final pass rate and win rate of trajectories, which raises questions to its reliability and interpretability.
@@ 2 ToolSandbox Design
At its core, ToolSandbox is a Python native LLM testing environment, with Execution Context as world state abstraction and Python functions as Tools, where User, Agent and Execution Environment communicate with each other through a Message Bus to complete a task, which is evaluated against predefined Milestones and Minefields. As shown in Figure 2, a typical test case starts with the User speaking to the Agent. From then on, the role being addressed gets to speak next, until the end state is reached. Upon receiving a User request, an Agent can decide to respond to the User asking for more information, or inform the Execution Environment to execute a Tool, providing intended tool name and arguments. The Execution Environment executes the Tool in an code.InteractiveConsole, Foundation (2024), which depending on the Tool modifies the world state stored in the Execution Context, and responds to the Agent. Once the User decides the task has been completed, it informs the Execution Environment to execute the end_conversation tool, which puts the system in the end state, ready to be evaluated based on the dialog’s similarity to Milestones and Minefields. A thorough walk-through of ToolSandbox implementation details can be found in Appendix A. The remainder of this section focuses on the key components that enables stateful, conversational and interactive evaluation.
2.1 Stateful
To construct challenging reasoning scenarios, ToolSandbox includes a set of carefully designed stateful tools, defined as tools that inspects, depends on or manipulates world states. These world states include:
Cellular service
Tools that require cellular service (e.g., send_message) depend on it to be true.
Wifi
RapidAPI tools (e.g., search_stock) depend on it to be true.
Location Service
All tools that utilize current location (e.g., get_current_location) depend on it to be true.
Low battery mode
All tools that turn on cellular service, wifi, and location service status depend on it to be false, creating nested state dependency.
Stateful tools cover 44% of ToolSandbox toolbox. Naturally these tools forms an implicit dependency between each other, testing the agent’s ability to maintain an internal image of world states and tool call stack. As an example, in Figure 1, when send_message tool is called while cellular service is off, a ConnectionError is raised, and the Agent should utilize set_cellular_service to resolve the error, understand that cellular service should be turned on now, and retry send_message.
2.2 Conversational
@@
On-policy conversational roll-out is supported by an LLM (GPT-4o) powered user simulator and carefully calibrated prompting design. The simulator represents a human interacting with an Agent, hoping to complete a task through possibly multiple rounds of conversation. When the User simulator decides the task has been completed, or could not be completed, it can terminate the conversation using the end_conversation tool, which is the single tool available to it. As related studies in user simulation Zhang et al. (2024); Sekulic et al. (2024) suggest, one should include the user’s overall goal in the simulator’s system prompt. However, we found this is often insufficient for the complex interactions in ToolSandbox, and can lead to two categories of failures. In some cases, it is infeasible for an LLM simulated user to judge task completion, or provide follow-up information with only access to the user goal, and not the expected result, which could lead to hallucination. Also, with only a single system prompt, the simulated user could be derailed by the tool-use agent, failing to follow instructions. Examples of these failures can be found in Appendix A.4.
@@
In light of this, we propose two additional components in user simulator prompts: Knowledge Boundary, which inform the user simulator what it should and should not know, providing partial access to expected result, combating hallucination. And Demonstration, which provides few shot example dialogs to the user simulator. Prompt examples can be found in Appendix A.4. Note that demonstration is only visible to the user simulator and not the agent. We performed an ablation study for these components in Table 2. With both approaches combined, the LLM simulated user achieves the lowest error rate in all categories. User simulator error rate is also found to be consistent across agents, shown in Table 3, which should not affect the agent accuracy comparison in Table 5.
@@
GPT-4o Claude-3-Opus Gemini-1.5-Pro
Hallucination 6.90±1.45 6.40±0.97 7.15±0.71
IF 1.11±0.84 1.38±0.69 0.92±0.49
Total Error 8.02±1.36 7.78±0.52 8.07±1.03
Table 3: Percentage of user simulation failures in each failure category for each agent model. Mean and std collected from 4 repeated trials each containing 1032 trajectories. Both Knowledge Boundary and Demonstration are applied.
2.3 Interactive
With an stateful, conversational and interactive environment, evaluation trajectories are highly dynamic. Multiple trajectories can lead to the same outcome. A given task may be completed using different tools, the same tools in a different order, or through trial and error, and the evaluation strategy has to be flexible enough to accommodate for that. To combat this, we developed an evaluation strategy based on Milestones and Minefields, which defines key events that must or must not happen in a trajectory, allowing us to evaluate any trajectory with rich intermediate and final execution signals, providing deeper understanding of the model performance. An example can be found in Figure 10.
@@ 2.3 Interactive
Each milestone also defines a similarity measure which calculates a 0 to 1 similarity between each turn and the milestone. Types of available similarity measures are introduced in Appendix A.7. Milestones form a directed acyclic graph (DAG) based on temporal dependency. To evaluate a trajectory against a milestone DAG, we find the highest averaged similarity $\text{score}_{M+}$ among all possible mappings between turns and milestones, given that the resulting chronological milestone sequence is a topological sort of the DAG. Task efficiency is not considered by Milestones, and is instead tracked by a complementary turn count metric shown in Appendix D.3. We introduce the milestone matching process with more details in Appendix A.7.
@@ 3 Test Scenarios
Avg Turns Avg Tool calls Test cases Tools
ToolSandbox 13.9 3.80 1032 34
BFCL 2.00 0.78 2000 1193
ToolEval 7.53 1.46 1625 3917
API-Bank 3.88 2.04 261 73
ToolTalk 7.42 3.68 78 28
$\tau$-bench 29.33 4.48 165 24
Table 4: Statistics between ToolSandbox and other tool-use benchmarks.
A test scenario is defined by the initial world state, the initial messages, the available tools and the evaluation milestone and minefields, as illustrated by the light blue boxes in Figure 1. ToolSandbox contains 1032 test scenarios meticulously crafted by 2 internal domain experts to capture challenging tool-use scenarios, with human authored and carefully calibrated Milestones and Minefields to support evaluation. One annotator is tasked to create test scenarios, while the other acts as an agent to validate milestones and minefields. We designed a rigorous annotation process to ensure coverage across realistic, complex use case scenarios, detailed in Appendix B.2. Statistics comparison between ToolSandbox and other benchmarks can be found in Table 4.
We designed 34 tools in ToolSandbox, covering 11 domains including Contact, Messaging, Reminder, System settings, Time utilities, Math utilities, Map, Weather, Stock, Conversion, and Holiday, backed by python native implementation when possible, carefully selected RapidAPI endpoints when necessary. Additional details about tool domain coverage and design principles can be found in Appendix B.3. Tools in ToolSandbox are designed to be representative, diverse and composable in conversational dialogs, while making tool count manageable for milestone annotation. As a result, ToolSandbox test scenarios contain on average much higher number of tool calls and turns per dialog compared to other benchmarks.
To closely inspect the intricate challenges in LLM tool-use applications, test scenarios are organized into the following categories:
@@
These categories apply to scenarios where one / multiple tool calls are needed to fulfill the user task. Examples are shown in Appendix C.2. Note that this definition is different from the Berkeley Function-Calling leaderboard Yan et al. (2024), which resembles distraction tools in ToolSandbox described in Tool Augmentation.
@@
Tool Augmentation
Orthogonal to the above categories, to support ablation studies on the effect of tool schema representation on agent accuracy, we have implemented multiple tool augmentations, including adding distraction tools, making tool or argument names less informative, and removing argument descriptions or type hints. For more details on the augmentations please refer to Appendix A.2.1.
Categories including Multiple Tool Call, Multiple User Turn, State Dependency and Insufficient Information are difficult challenges which requires complex reasoning capability from the agent, 85% of ToolSandbox scenarios are associated with at least one of these challenging categories. Further scenario statistics including category wise breakdown and milestone coverage can be found in Appendix B.4.
@@ 4 Evaluation Results
When evaluating the models, all models use the same minimalist prompt shown in Figure 8 for comparison fairness. We do not include additional prompt engineering for the models, as we consider prompt engineering gains orthogonal to the innate model capability surfaced by simpler prompting. Table 5 shows the average similarity for each of the scenario categories described in Section 3. Additional prompting experiments can be found in Appendix 9.
@@
There is a significant performance gap between proprietary and open source models, with the best performing open source model Hermes interstellarninja et al. lagging more than 20 points behind the second to last proprietary model Claude-3-Haiku Anthropic (2024). This is partly due to the fact that models like Gorilla Patil et al. (2023) and Command-R Cohere and for AI (2024) are incapable of consuming tool responses, as shown in Appendix D.2. They can theoretically solve Single Tool Call test scenarios, but would fail in any scenario that requires multiple tool calls. As for Hermes and Mistral Jiang et al. (2023), both models struggle at identifying when a tool call should be issued. Mistral for example would often mistake a tool-use scenario for a code generation task, as shown in Figure 11. These models’ subpar performance unexpectedly caused them to achieve higher rating in the Insufficient Information category, which rewards the model for not generating hallucinated tool calls or arguments when provided tools are insufficient to complete the task. This should be considered a side effect instead of a positive outcome.
@@
The State Dependency category shows an interesting trend where, larger models like GPT-4 Achiam et al. (2023) and Claude-3-Opus perform significantly worse than mid to smaller sized models like GPT-3.5-Turbo and Claude-3-Sonnet. This is due to erroneous parallel tool calls in face of state dependency. As mentioned in Appendix A.6, the Execution Environment always surfaces race conditions when present. Larger models like GPT-4 and Claude-3-Opus are prune to issuing parallel tool calls even for dependent tools, leading to a performance deficiency. An example is shown in Figure 17. Nested state dependency is also tricky to solve efficiently. As shown in Figure 18, models often forget about open issues and would not optimally backtrack, leading to repeated errors and as a result a much higher than optimal turn count.
@@
Robustness against tool augmentations seems to vary model by model. While adding distraction tools, Claude-3-Sonnet drops almost 10 points between 0 distraction tools, and making all ToolSandbox tools available. GPT-4o is particularly susceptible to Tool Description Scrambling, GPT-4 pays extra attention to argument descriptions, and Gemini-1.5 doesn’t do well with Argument Type Scrambling.